# Hanle Effect for Lifetime Analysis: Li-like Ions

### v1

@@ 1 Introduction
The resonant scattering in the above mentioned experiment took place via the transition from the $1\mathrm{s}^{2}$ ${}^{1}\mathrm{S}_{0}$ ground state to one of the $1\mathrm{s}n\mathrm{p}$ ${}^{1}\mathrm{P}_{1}$ excited states. The present work extends the theoretical investigation of lifetime measurements based on the Hanle effect to $J_{i}=1/2\to J\nu=1/2$ and $J_{i}=1/2\to J\nu=3/2$ transitions, as can be observed, for instance, in lithium-like ions. Our analysis starts in Sec. 2 with establishing the appropriate scattering geometry, based on the experimental setup described in Ref. [4]. The theoretical description of the scattering process, which is relying on a second-order perturbation approach, is given then in Sec. 3. Here, in particular, we discuss the angle-differential cross section for the scattering of linearly polarized incident light. In Sec. 4, we show that the ratio of cross sections for outgoing photons emitted either in parallel or perpendicular direction to the polarization vector of incident ones is very sensitive both to the external magnetic field and the lifetime of an intermediate $J_{\nu}=3/2$ state. Lifetime measurements of the excited states of Li-like ions can be performed by employing this sensitivity for the range $10^{-13}$ s $\leq\tau_{\nu}\leq 10^{-10}$ s. A summary of these results and their potential for future experimental applications are discussed in Sec. 5. Relativistic units ($\hbar=c=m_{e}=1$) are used throughout this paper unless stated otherwise.
@@ 4 Results
We will investigate how the ratio of these cross sections, $\sigma_{\parallel}/\sigma_{\perp}$, depends on the external magnetic field and lifetime of an excited ionic state.
@@ 4.1 Scattering via a $J_{\nu}=1/2$ state
In this subsection, we briefly discuss a $J_{i}=1/2\to J_{\nu}=1/2\to J_{i}=1/2$ transition, such as for example the photon scattering by Li-like ions in the $1\mathrm{s}^{2}2\mathrm{s}$ ${}^{2}\mathrm{S}_{1/2}$ ground state via an intermediate $1\mathrm{s}^{2}n\mathrm{p}$ ${}^{2}\mathrm{P}_{1/2}$ level. Using Eqs. (3.1)-(3.4), one finds that for $J_{i}=J_{\nu}=1/2$, the angular distribution of scattered photons in the field-free case, $B=0$, is isotropic [6], resulting in the ratio $\sigma_{\parallel}/\sigma_{\perp}=1$. This isotropy persists even in the presence of an external magnetic field when accounting for the modified energies from Eq. (3.3). Consequently, the ratio remains $\sigma_{\parallel}/\sigma_{\perp}=1$, making it impossible to extract the lifetimes of excited states from the analysis of the differential cross section.
@@ 4.2 Scattering via a $J_{\nu}=3/2$ state
thus leading to the ratio $\sigma_{\parallel}/\sigma_{\perp}=2/5$. When a magnetic field is applied, the cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$ and, hence, the ratio $\sigma_{\parallel}/\sigma_{\perp}$ are modified. While the exact analytical expression for the cross section ratio in the presence of an external magnetic field is cumbersome, and therefore is not presented here, we investigate it numerically. To perform this quantitative analysis, we focus below on the resonant $1\mathrm{s}^{2}2\mathrm{s}$ ${}^{2}\mathrm{S}_{1/2}\to 1\mathrm{s}^{2}n\mathrm{p}$ ${}^{2}\mathrm{P}_{3/2}$ scattering. As seen from Eq. (3.3), the computation of the Zeeman splitting, and hence the $g$-factors, of the involved states is needed. For low-$Z$ ions, the $g$-factors can be approximated as $g_{{}^{2}\mathrm{S}_{1/2}}=2$ and $g_{{}^{2}\mathrm{P}_{3/2}}=4/3$. Moreover, to account for a realistic experimental scenario of a finite frequency width of the incoming radiation, we average the cross section (3.4) over a Gaussian distribution:
@@
The ratio $\tilde{\sigma}_{\parallel}/\tilde{\sigma}_{\perp}$ of the frequency averaged cross sections depends both on the strength of the external magnetic field and the natural width $\Gamma_{\nu}$ of an excited state, directly related to its lifetime $\tau_{\nu}=\hbar/\Gamma_{\nu}$. In order to better understand this sensitivity, we display in Fig. 2 the ratio $\tilde{\sigma}_{\parallel}/\tilde{\sigma}_{\perp}$ calculated for different magnetic field strengths, $B=0.2$ T (blue dashed line), $B=0.85$ T (red solid line), $B=6$ T (green dotted line), and as a function of the lifetime $\tau_{\nu}$. Hence, we formally use $\tau_{\nu}$ as a free parameter of our calculations with the goal to specify a feasible range for lifetime measurements. For reference, by the vertical lines in Fig. 2 we mark as examples actual lifetimes of $1\mathrm{s}^{2}3\mathrm{p}$ ${}^{2}\mathrm{P}_{3/2}$ states in Li-like F, Mg, and S ions.
@@
Figure 2: The ratio of frequency averaged cross sections for photons scattered within ($\tilde{\sigma}_{\parallel}$) or perpendicular ($\tilde{\sigma}_{\perp}$) to the polarization plane of the incident radiation as a function of the lifetime of the excited state. The calculations were performed for external magnetic field strengths of $B=0.2$  T (blue dashed line), $B=0.85$ T (red solid line), and $B=6$ T (green dotted line). The vertical line present the lifetimes of the $1\mathrm{s}^{2}3\mathrm{p}$ ${}^{2}\mathrm{P}_{3/2}$ state in Li-like F, Mg and S ions.

### v2

@@ 1 Introduction
The resonant scattering in the above mentioned experiment took place via the transition from the $1\mathrm{s}^{2}$ ${}^{1}\mathrm{S}_{0}$ ground state to one of the $1\mathrm{s}n\mathrm{p}$ ${}^{1}\mathrm{P}_{1}$ excited states. The present work extends the theoretical investigation of lifetime measurements based on the Hanle effect to the scattering of photons by ions in the ground state with angular momentum $J_{i}=1/2$ via excited states with angular momentum $J_{\nu}=1/2$ or $J_{\nu}=3/2$, as it can be observed, for instance, in Li-like ions. Our analysis starts in Sec. 2 with establishing the appropriate scattering geometry, based on the experimental setup described in Ref. [4]. The theoretical description of the scattering process, which is relying on a second-order perturbation approach, is given then in Sec. 3. Here, in particular, we discuss the angle-differential cross section for the scattering of linearly polarized incident light. In Sec. 4, we show that the ratio of cross sections for outgoing photons emitted either in parallel or perpendicular direction to the polarization vector of incident ones is very sensitive both to the external magnetic field and the lifetime of an intermediate $J_{\nu}=3/2$ state. Lifetime measurements of the excited states of Li-like ions can be performed by employing this sensitivity for the range $10^{-13}$ s $\leq\tau_{\nu}\leq 10^{-10}$ s. A summary of these results and their potential for future experimental applications are discussed in Sec. 5. Relativistic units ($\hbar=c=m_{e}=1$) are used throughout this paper unless stated otherwise.
@@ 4 Results
We will investigate how the ratio of these cross sections, $\sigma_{\parallel}/\sigma_{\perp}$, depends on the external magnetic field and lifetime of an excited ionic state. Moreover, we neglect any effects of hyperfine structure on the scattering process, by restricting our analysis to ions with nuclear spin $I=0$.
@@ 4.1 Scattering via a $J_{\nu}=1/2$ state
In this subsection, we briefly discuss a $J_{i}=1/2\to J_{\nu}=1/2\to J_{i}=1/2$ transition, such as for example the photon scattering by Li-like ions in the $1\mathrm{s}^{2}2\mathrm{s}$ ${}^{2}\mathrm{S}_{1/2}$ ground state via an intermediate $1\mathrm{s}^{2}n\mathrm{p}$ ${}^{2}\mathrm{P}_{1/2}$ level. Using Eqs. (3.1)-(3.4), one finds that for $J_{i}=J_{\nu}=1/2$, the angular distribution of scattered photons in the field-free case, $B=0$, is isotropic [6], resulting in the ratio $\sigma_{\parallel}/\sigma_{\perp}=1$. This isotropy persists even in the presence of an external magnetic field when accounting for the modified energies from Eq. (3.3). Consequently, the ratio remains $\sigma_{\parallel}/\sigma_{\perp}=1$, making it impossible to extract the lifetimes of excited states from the analysis of the angular distribution of emitted photons.
One should note that a different experimental setup was proposed in Refs. [16, 17], which allows lifetime determination utilizing the Hanle effect also for $J_{i}=1/2\to J_{\nu}=1/2\to J_{i}=1/2$ transitions. This setup relies on the scattering of circularly polarized incoming light as well as on the detection of circularly polarized outgoing photons. Moreover, a scan over the external magnetic field strength is performed to observe the zero-field level crossing of the degenerate magnetic sublevels of the excited state. This scenario, requires the production and detection of circularly polarized radiation, which is well established in the visible light domain, but is more challenging in the EUV and X-ray regime.
@@ 4.2 Scattering via a $J_{\nu}=3/2$ state
thus leading to the ratio $\sigma_{\parallel}/\sigma_{\perp}=2/5$. When a magnetic field is applied, the cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$ and, hence, the ratio $\sigma_{\parallel}/\sigma_{\perp}$ is modified. While the exact analytical expression for the cross section ratio in the presence of an external magnetic field is cumbersome, and therefore is not presented here, we investigate it numerically. To perform this quantitative analysis, we focus below on the resonant $1\mathrm{s}^{2}2\mathrm{s}$ ${}^{2}\mathrm{S}_{1/2}\to 1\mathrm{s}^{2}n\mathrm{p}$ ${}^{2}\mathrm{P}_{3/2}$ scattering. As seen from Eq. (3.3), the computation of the Zeeman splitting, and hence the $g$-factors, of the involved states is needed. For low-$Z$ ions, the $g$-factors can be approximated as $g_{{}^{2}\mathrm{S}_{1/2}}=2$ and $g_{{}^{2}\mathrm{P}_{3/2}}=4/3$. Moreover, to account for a realistic experimental scenario of a finite frequency width of the incoming radiation, we average the cross section (3.4) over a Gaussian distribution:
@@
The ratio $\tilde{\sigma}_{\parallel}/\tilde{\sigma}_{\perp}$ of the frequency averaged cross sections depends both on the strength of the external magnetic field and the natural width $\Gamma_{\nu}$ of an excited state, directly related to its lifetime $\tau_{\nu}=\hbar/\Gamma_{\nu}$. In order to better understand this sensitivity, we display in Fig. 2 the ratio $\tilde{\sigma}_{\parallel}/\tilde{\sigma}_{\perp}$ calculated for different magnetic field strengths, $B=0.2$ T (blue dashed line), $B=0.85$ T (red solid line), $B=6$ T (green dotted line), and as a function of the lifetime $\tau_{\nu}$. Hence, we formally use $\tau_{\nu}$ as a free parameter of our calculations with the goal to specify a feasible range for lifetime measurements. For reference, by the vertical lines in Fig. 2 we mark as examples actual lifetimes of $1\mathrm{s}^{2}3\mathrm{p}$ ${}^{2}\mathrm{P}_{3/2}$ states in Li-like O, Mg, and S ions.
@@
Figure 2: The ratio of frequency averaged cross sections for photons scattered within ($\tilde{\sigma}_{\parallel}$) or perpendicular ($\tilde{\sigma}_{\perp}$) to the polarization plane of the incident radiation as a function of the lifetime of the excited state. The calculations were performed for external magnetic field strengths of $B=0.2$  T (blue dashed line), $B=0.85$ T (red solid line), and $B=6$ T (green dotted line). The vertical line present the lifetimes of the $1\mathrm{s}^{2}3\mathrm{p}$ ${}^{2}\mathrm{P}_{3/2}$ state in Li-like O, Mg and S ions.